Reports of 30757
VIS (EX)
takahiro.yamamoto - 19:32 Friday 04 April 2025 (33253) Print this report
Comment to Installation of Coil Driver Switcher (33152)
[Dan, YamaT]

Abstract

We took back the BIO cables (CB{I,O}_L32_C1) for the BFV coil driver from the coil driver switcher,
and stole again the BIO cables (CB{I,O}_L32_C0) for the coil driver switcher from the PI coil driver.
After then, only one BO cable (CBO_H32_C1) is unused, so it connected to the CTRL port of the PI coil driver.
As a result, there is no readback of de-whitening filters for PI though it can be switched.
After our work, we engaged two stages of de-whitening filters for PI with confirming it by a sounds of relay switch.
 

Details

In the previous work klog#33152, we stole the BIO cables (CB{I,O}_L32_C1) for the coil driver switcher from the BFV coil driver. After then, Kenta and Ushiba-kun found the DAC noise contamination of GAS and BF stages (klog#33197). At that time they couldn't enable de-whitening filters for BFV because of the stolen BIO cables. As the temporal solution for this issue, Ushiba-kun inserted a config board for the BFV coil driver (klog#33223). In order to remove this temporal solution, we rewired the BIO cables.

Trying installation of a new BIO cards
At first, we tried to install a new BIO card as the 6th one on the IO chassis in the K1EX1 rack. But real-time front-end couldn't find any PCIe cards on the IO chassis when the 6th BIO cards were installed. According to my experience in the test stand, we was able to use 6 or 7 BIO cards on one IO chassis. Ambiguity of a number of cards seemed to come from an used cable between real-time front-end and IO chassis, the power supply situation, etc.

Though I'm not sure an exact reason why we reached a limit of a number of BIO cards as only 5, I guessed that it's also related to the total number of cards including also ADC, DAC, and BO. I was able to use 6 or 7 BIO cards with 1 ADC and 1DAC cards on the test stand. On the other hand, 3 ADC, 3 DAC, and 1 BO cards are installed on K1EX1. Anyway, we couldn't find a way to use the 6th BIO cards on K1EX1 though we tried to reboot some times and to use another PCIe slots.

Rewiring BIO cables
Because we found that it was impossible to use the 6th BIO card, we changed our strategy to the next best way as that BIO cables for the coil driver switcher is stolen from the PI coil driver instead of the BFV coil driver (DAC noise contamination from PI should be a smaller impact than one from BFV). So we removed the config board on BFV coil driver added by Ushiba-kun and took back the BIO cables from the coil driver switcher to BFV coil driver. Then we stole BIO cables again from PI coil driver in order to use on the coil driver switcher.

After rewiring, we found only one BO cable which was not used. A paired BI cable with this unused BO cable is apparently used for release sensors and the release sensors has only monitor channels. So we used this BO cable for PI coil driver for a control of de-whitening filters. Because there is no enough BI cables for PI coil driver, we cannot check the readback signal to ensure the state of filter engagement. So we confirmed to engage two stages of de-whiteninig filters by hearing a sound by relay switch. In common case, a state of the de-whitening filters for PI is never touched. So it might be enough to use a config board. If we obtained a benefit of using BO even an absence of BI, it's a time when we need to do a health check for ETMX. It's difficult to do the health check with enabling de-whitening filters because of DAC saturation.

Operation check
After the installation, we confirmed that de-whitening filter for all stages can be switched one-by-one and the test mass can be driven with LPCD. Fig.1 and Fig.2 shows the TM oplev signals when excitation was enabled from HPCD path and LPCD path, respectively. (I manually stopped two ndscope window for oplev signal and DAC output, so time epoch of two window has slightly a bias. Sorry for the lack of clarity.) We can see some excess on the oplev signals, so the coil driver output seemed to work fine. After then, we requested ALS_DARM and confirmed that changes in this work don't affect the lock acquisition.
Images attached to this comment
PEM (Center)
tatsuki.washimi - 18:29 Friday 04 April 2025 (33259) Print this report
Comment to ACCs & Shakers for the PR3 chamber (33254)

I checked these ACCs' signals. The ACC on the leg ("BNC1" channel) is quite smaller than others.
It looks like it is not working, but it also responded when I jumped nearby.

Images attached to this comment
PEM (Center)
tatsuki.washimi - 18:05 Friday 04 April 2025 (33258) Print this report
Comment to Back the MCF Seismometer to near the MCF chamber (33252)

Its spectrum looks not strange

Images attached to this comment
PEM (Center)
tatsuki.washimi - 18:01 Friday 04 April 2025 (33257) Print this report
Comment to PEM sensors around the OMC (33250)

their spectrums are checked. not strange.

Images attached to this comment
PEM (Center)
tatsuki.washimi - 17:40 Friday 04 April 2025 (33254) Print this report
ACCs & Shakers for the PR3 chamber

We located 4 accelerometers and 2 shakers for the PR3 chamber.

Images attached to this report
Comments to this report:
tatsuki.washimi - 18:29 Friday 04 April 2025 (33259) Print this report

I checked these ACCs' signals. The ACC on the leg ("BNC1" channel) is quite smaller than others.
It looks like it is not working, but it also responded when I jumped nearby.

Images attached to this comment
VAC (Valves & Pumps)
koji.nakagaki - 17:15 Friday 04 April 2025 (33251) Print this report
Comment to Automatic closure installed on the IFI large gate valve (32852)

[Yasui, Nakagaki]

We set a threshold on the IFI gate valve auto-closure device and began monitoring.

The vacuum gauge threshold was set at 5.4e-5[Pa]. 

Images attached to this comment
PEM (EX)
takaaki.yokozawa - 17:10 Friday 04 April 2025 (33256) Print this report
Comment to Checking of the EXV Seiemometer cablings (33249)
During this work, we noticed that the 1 Hz triangle like signal detected from EXV seismometer.
From Washimi-san's investigation, this would be caused by the mass mis-centering.
So we performed the mass centering.

I compared the spectrum around noon (red) and early morning (blue)
Images attached to this comment
DetChar (General)
Takahiro S. Yamamoto - 17:10 Friday 04 April 2025 (33255) Print this report
How far can KAGRA observe BBHs with the current sensitivity?

(TSYamamoto, Yuzurihara)

We estimated BBH's inspiral range with the current sensitivity. We used gwpy.astro.inspiral_range with gwpy=3.0.9. The result is summarized here.

PEM (Center)
tatsuki.washimi - 17:00 Friday 04 April 2025 (33252) Print this report
Back the MCF Seismometer to near the MCF chamber

I backed the MCF Seismometer, used near the IFI chamber, to near the MCF chamber.

Images attached to this report
Comments to this report:
tatsuki.washimi - 18:05 Friday 04 April 2025 (33258) Print this report

Its spectrum looks not strange

Images attached to this comment
PEM (Center)
tatsuki.washimi - 16:56 Friday 04 April 2025 (33250) Print this report
PEM sensors around the OMC

I reconnected the PEM sensors around the OMC area.

  • Trillium compact seismometer on the ground
  • Ground voltage probe for the chamber
  • Microphone in the booth
  • Accelerometer on a bellows leg
Images attached to this report
Comments to this report:
tatsuki.washimi - 18:01 Friday 04 April 2025 (33257) Print this report

their spectrums are checked. not strange.

Images attached to this comment
PEM (EX)
tatsuki.washimi - 16:43 Friday 04 April 2025 (33249) Print this report
Checking of the EXV Seiemometer cablings

[YokozaWashimi]

We checked the cabling status of the EXV Seiemometer (Trillium120QA, S1707551)

  • The whiteninf filter is 15-pin type (S1605077), but a strange cable (the rack side is a D-sub 15-pin with 9 pins, and the seismometer side is a D-sub 9-pin.
  • A DC power tap is located inside a plastic box.
  • DC power to operate the Trillium was provided by a D-sub 3-pin cable with 2 pins.
  • The connection between the  Trillium cable and the D-sub 9/3 cables is a handmade one. It looks very weak. The scrues are not tightened.
Images attached to this report
Comments to this report:
takaaki.yokozawa - 17:10 Friday 04 April 2025 (33256) Print this report
During this work, we noticed that the 1 Hz triangle like signal detected from EXV seismometer.
From Washimi-san's investigation, this would be caused by the mass mis-centering.
So we performed the mass centering.

I compared the spectrum around noon (red) and early morning (blue)
Images attached to this comment
CAL (General)
hirotaka.yuzurihara - 16:41 Friday 04 April 2025 (33240) Print this report
TCam photo session 20250404
PEM (Center)
tatsuki.washimi - 16:21 Friday 04 April 2025 (33248) Print this report
Checking of the IXV Seiemometer cablings

[YokozaWashimi]

We checked the cabling status of the  IXV Seiemometer (Trillium120QA, S1707552)

  • The whiteninf filter is 15-pin type (S1605080), but conberted to a D-sub 9-pin.
  • In the plastic box, a barometer (S/N L4420370) was located.
  • DC power to operate the Trillium was provided by a D-sub 3-pin cable.
  • The connection between the  Trillium cable and the D-sub 9/3 cables is a handmade one. It looks very weak. The scrues are not tightened.

These cables can be replaced with one D-sub 15 cable.

Images attached to this report
VIS (EX)
satoru.ikeda - 15:51 Friday 04 April 2025 (33247) Print this report
Comment to Installation of Coil Driver Switcher (33152)

YamaT-san, DanChen-san, Ushiba-san, Ikeda

Overview

Work related to K-Log#33152 : Installation of Coil Driver Switcher

The model of k1ex1 was modefied for BIO connection of Low Power Coil Driver.
YamaT-san and others reported that they could not recognize all cards when the sixth DIO6464 card was inserted, so the existing BIO connection was changed.

Details of the change

BIO connection changes
 Before change
  k1visetmxlsc
   IMB: H32_C3
   TP: L32_C4
  k1visetmxt
   PI: L32_C0

 After change
  k1visetmxlsc
   IMB: remove
   TP (LPCD): H32_C3 (Moved from IMB) <- Reuse the entire coil driver from IMB, so use the wiring as IMB
   CoilDriver_SW: L32_C0 (Move from PI)
   TM_HP: L32_C4 (Move from TP)
  k1visetmxt
    PI: H32_C4 (BO: BO side of Release_Sensor is unused, so use this, BI:remove)

Model change point
PAYLOAD_MASTER/
  Add the following new block
 CRYO_PAYLOAD_16K_HP_MASTER
  Copy from CRYO_PAYLOAD_16K_MASTER and change BIO,TM block
  PAYLOAD_OPLEV_HP_TYPE_A_16K
    Copy from PAYLOAD_OPLEV_TYPE_A_16K and change MASTER, add COILOUTF_HP
  PAYLOADMASTER_HP
  PAYLOADBIO_HP
  A_PAYLOADBIO_HP_ENCODE  
  A_PAYLOADBIO_HP_DECODE
   Copy the original and add HP-related

 CRYO_PAYLOAD_16K_HP_MASTER
  Replaced with new block

k1visetmxt:
  Only change BIO connection

k1visetmxlsc: 
  Changed from CRYO_PAYLOAD_16K_MASTER to CRYO_PAYLOAD_16K_HP_MASTER
  Change of BIO connection
  DAC_1 In12-15: Change from TM_{H1,H2,H3,H4} to TM_{H1,H2,H3,H4}_HP
  DAC_2 In12-15: Change from GND to TM_{H1,H2,H3,H4}
  ETMX block changed from CRYO_PAYLOAD_16K_MASTER to CRYO_PAYLOAD_16K_HP_MASTER

SDF overwrote the EPICS value of TM_COILOUTF to safe.snap after renaming from TM_COILOUTF to TM_COILOUTF_HP.
Filter banks were copied from TM_COILOUTF to TM_COILOUTF_HP.
 

Images attached to this comment
Non-image files attached to this comment
CRY (Cryo-payload EY)
shinji.miyoki - 13:50 Friday 04 April 2025 (33246) Print this report
Comment to Apply current to the IM heater on EY payload (28420)

The current was decreased by changing the voltage to 9.500V.

Images attached to this comment
DetChar (General)
hirotaka.yuzurihara - 13:34 Friday 04 April 2025 (33245) Print this report
Comment to Added new SummaryPages states (33241)

> Locked(IFO)
> The state is when the IFO guardian is 100.
Supplemental comment: the condition to be K1:GRD-IFO_STATE_N == 100 is `K1:GRD-LSC_LOCK_REQUEST_N > 1400`, which is described in JGWdoc.

DetChar (General)
shoichi.oshino - 11:49 Friday 04 April 2025 (33244) Print this report
modified plotting range of PSL table temperature and PSL QPD
I modified plotting range of PSL table temperature and PSL QPD graph on SummaryPages.
DetChar (General)
shoichi.oshino - 11:28 Friday 04 April 2025 (33243) Print this report
Comment to Omicron stopped at 4/2 (33212)
Due to this trouble, Omicron didn't work at 4/2.
Therefore, I reran Omicron on 4/2 and in the morning on 4/3.
After finishing Omicron process, I also reran Hveto process.
DetChar (General)
shoichi.oshino - 11:26 Friday 04 April 2025 (33242) Print this report
Comment to Added new SummaryPages states (33241)
We also added "Locked (RF)" state that shows the interferometer condition at PRFPMI RF state.
The state is when the LSC_LOCK guardian is 1301.
DetChar (General)
shoichi.oshino - 10:49 Friday 04 April 2025 (33241) Print this report
Added new SummaryPages states
[Ushiba, Yuzurihara, Oshino]

As discussed in the morning meeting, we added new SummaryPages states.

* Working
This state shows the interferometer condition at work during the day.
The state is when the LSC_LOCK guardian is 9998 or higher.

* Silent run
This state shows the interferometer condition is stable.
The state is when the LSC_LOCK guardian is 9990.

Removed the Locked (LSC) state that is not currently in use.

The following states remain as they were before.

* Observing
This state shows that the interferometer is in the observation state.
The state is when the IFO guardian is 1000.

* Locked(IFO)
This state shows that the interferometer is in the OMC DC readout or later state.
The state is when the IFO guardian is 100.
Comments to this report:
shoichi.oshino - 11:26 Friday 04 April 2025 (33242) Print this report
We also added "Locked (RF)" state that shows the interferometer condition at PRFPMI RF state.
The state is when the LSC_LOCK guardian is 1301.
hirotaka.yuzurihara - 13:34 Friday 04 April 2025 (33245) Print this report

> Locked(IFO)
> The state is when the IFO guardian is 100.
Supplemental comment: the condition to be K1:GRD-IFO_STATE_N == 100 is `K1:GRD-LSC_LOCK_REQUEST_N > 1400`, which is described in JGWdoc.

DetChar (General)
hirotaka.yuzurihara - 9:18 Friday 04 April 2025 (33237) Print this report
Summary of recent segment for IPC glitch (2025/03~04/03)

As Yokozawa-san pointed in klog33232 and the control room members already recognized, recently the IPC glitches occurring frequently.

  • 3/21 n=6
  • 3/22 n=0
  • 3/23 n=2
  • 3/24 n=90
  • 3/25 n=74
  • 3/26~27 n=0
  • 3/28 n=50
  • 3/29 n=92
  • 3/30~31 n=0
  • 4/1 n=99
  • 4/2 n=66
  • 4/3 n=39

During O4a, we also saw many IPC glitches around noon and they concentrated in specific one hour. But, the recent IPC glitches occurred not in specific one hour and occurred in broad time. Roughly speaking the frequency is once per 20 minutes or shorter than 20 minutes. I attach the list of occurance times in JST. I didn't find the characteristic features in this list.

Non-image files attached to this report
MIF (General)
takaaki.yokozawa - 8:59 Friday 04 April 2025 (33239) Print this report
Comment to Measurement of the resonant peaks 45 Hz pitch in Type-A suspension (33236)
I generated the NB filter for the IX (45 Hz) using P04
I will check the performance of this filter after locking the interferometer
Images attached to this comment
MIF (General)
takaaki.yokozawa - 8:55 Friday 04 April 2025 (33236) Print this report
Measurement of the resonant peaks 45 Hz pitch in Type-A suspension
I tried to identify which suspension had the resonant peak of 45 Hz in pitch for the Type-A suspensions.
I injected 40 - 50 Hz white noise to MNP test filter bank. This is the same method with klog32983

Fig.1. IX about 45 Hz
Fig.2. IY about 44 Hz
Fig.3. EX about 44.5 Hz
Fig.4. EY about 44.4 Hz (cannot be seen in silent case)
Images attached to this report
Comments to this report:
takaaki.yokozawa - 8:59 Friday 04 April 2025 (33239) Print this report
I generated the NB filter for the IX (45 Hz) using P04
I will check the performance of this filter after locking the interferometer
Images attached to this comment
DetChar (General)
hirotaka.yuzurihara - 8:54 Friday 04 April 2025 (33235) Print this report
Comment to No scattered light noise and transient noise in PRCL feedback of 4/2 and 4/3 (33217)

This might be a jumping to conclusion... In the yesterday's data (21~22 JST), there was the scattered light noise in MICH feedback channel, even though there is notthing special in the PRCL feedback channel. Figure 1 and 2 show the example. Anyway, the situation for the scattered light noise is changing from the last month (before fixing the STM2).

Images attached to this comment
MIF (General)
takaaki.yokozawa - 7:43 Friday 04 April 2025 (33232) Print this report
Alignment work 250404
After the large earthquake, the alignment became worse and LSC_LOCK guardian stopped at the PRMI lock trial with ALS_DARM,CARM lock.
So, I performed the initial alignment only for the PRMI, then PRFPMI_RF_LOCKED succeeded.
Next, I tried to perform the OMC alignment with the following klog
https://klog.icrr.u-tokyo.ac.jp/osl/?r=33214
Then, interferometer became stable.

Today, IPC glitch like signal detected so much (once per several minutes).
Search Help
×

Warning

×